OpenStudy (anonymous):

Find the two square roots for the complex number. Answer in standard form. -2+2i√3

OpenStudy (ash2326):

@Ala123 Can you convert this into polar form?

OpenStudy (anonymous):

you mean Trig form ?

OpenStudy (ash2326):

The form \[re^{i \theta}\]

OpenStudy (anonymous):

what form is that? i only learned trig form for this

OpenStudy (ash2326):

Ok, write that down

OpenStudy (ash2326):

Yeah, just remembered you can use that also

OpenStudy (anonymous):

idk how to do that :|

OpenStudy (anonymous):

is that re^itheta ?

OpenStudy (ash2326):

You write that trigonometric form. We'll work with that

OpenStudy (anonymous):

(a+i*b)²=a²-b² +i*2*a*b=-2+2i√3 compare real and imaginary, recall a,b element from R

OpenStudy (anonymous):

ok 1 step at a time please -.-

OpenStudy (anonymous):

i know if you convert to trig form you gotta find r first

OpenStudy (anonymous):

r=√a^2+b^2

OpenStudy (anonymous):

in this case a^2 =-2 but what is b^2 ?

OpenStudy (anonymous):

b=2√3i ?

OpenStudy (ash2326):

I'll help you with that

OpenStudy (ash2326):

a=-2 and \(b=2\sqrt 3\) Don't include i

OpenStudy (anonymous):

@ash2326 if you put the 2√3 and you square it mean the root cancel out right ?

OpenStudy (ash2326):

Yeah \[ (2 \sqrt 3)^2= 2 \times 2 \times \sqrt 3\times \sqrt 3\]

OpenStudy (anonymous):

is it 6 ?

OpenStudy (anonymous):

oh it's 7

OpenStudy (ash2326):

Is this r?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

oh no i mean r=root 7

OpenStudy (anonymous):

root 11

OpenStudy (anonymous):

r=root 11

OpenStudy (ash2326):

How did you get that? \[r=\sqrt {2^2+ (2\sqrt 3)^2}=\sqrt {4+12}=4\]

OpenStudy (anonymous):

r=4

OpenStudy (anonymous):

there lol lol =.=

OpenStudy (ash2326):

ok, find the theta?

OpenStudy (anonymous):

theta can you just put tan inverse(b/a)

OpenStudy (anonymous):

i got theta=60 :D

OpenStudy (ash2326):

It's better to plot this first. If you plot this point then you'll know better what's theta?

OpenStudy (anonymous):

120 ?

OpenStudy (ash2326):

Yay:D that's right, now what's the r e^(itheta form)

OpenStudy (anonymous):

i only know Trig form =.= which is 4(cos120+ i sin120)

OpenStudy (anonymous):



OpenStudy (ash2326):

I made a mistake sorry:( Theta is not 120 Thanks @satellite73 for pointing out

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

@ash2326 no you are right, \(\theta =120\) if you are working in degrees.

OpenStudy (anonymous):

yeah, lol u just gave me a mini heart attack :O

OpenStudy (anonymous):

take the square root of 4, take half of 120 get \[2(\cos(60)+i\sin(60))\] evaluate

OpenStudy (anonymous):

then go half way across the circle

OpenStudy (anonymous):

2cis60 and 2cis240 and idk how to turn into standard form :D

OpenStudy (mertsj):

\[2(\cos60+isin60)=2(\frac{1}{2}+i\frac{\sqrt{3}}{2})=1+\sqrt{3}i\]

OpenStudy (anonymous):

oh yeah :D

OpenStudy (mertsj):

\[2(\cos 240+i \sin 240)=2(\frac{-1}{2}-\frac{\sqrt{3}}{2})=-1-\sqrt{3}\]

OpenStudy (mertsj):

i. I forgot the i

OpenStudy (anonymous):

ok i'll include it thanks Mertsj

OpenStudy (mertsj):

yw